Classify each of the following as true or false:
(a) If a number is divisible by 6, then it is divisible by 2 and 3.
(b) If a number is divisible by 2 and 3, then it is divisible by 6.
(c) If a number is divisible by 2 and 4, then it is divisible by 8.
(d) If a number is divisible by 8, then it is divisible by 2 and 4.

Answer by richwmiller(17219)

a) true 12/3=4 12/2=6 12/6=2
b) true 6,12,18 6/6=1 12/6=2 18/6=3
c) false 12/2=6 12/4=3 12/8=3/2
d) true
8/8=1 8/4=2 8/2=4
16/8=2 16/2=8 16/4=4